Grade 10Electric Current

Truck A travels at a constant speed of 50 miles/h on a straight road. Truck B travels at a constant speed of 70 miles/h in the same direction. Truck A passes a gas station at 12 noon. Truck B passes the same gas station at 12:30 pm. At what time does truck B catch up with truck A? How far are the trucks from the gas station when this happens?

Aditi Chauhan
12 Years agoGrade 10

1 Answer

Ram Kushwah
5 Years ago
At the gas station the time difference between both trucks is 30 minutes
So when truck B reaches at gas station ,mean while Truck has passed
distance=1/2*50=25 miles
Now the speed difference =70-50=20 miles/hr
Thus the truck B will compensate these 25 miles
in=25/20=1.25 hr
Hence the truck B will catch truck B after 1.25 hrs from 12:300 Pm
=12:30+1.25 hr=12:30+1hr 15 min
=1:45 PM
Thus trucks will meet at 1:45 hrs
****************************************************
II Distance of trucks from gas station  when they meet
=Distance travelled by truck B in 1.25 hrs
=1.25*70=87.50 miles from gas station